Find the equation of the curve having y’ = 2x – 5 that passes through (5,4). 


1
Expert's answer
2022-04-25T17:48:25-0400

Let us find the equation of the curve having "y' = 2x \u2013 5" that passes through "(5,4)." It follows that "y=x^2-5x+C," and hence "4=5^2-5^2+C=C." We conclude that the equation of the curve is "y=x^2-5x+4."